OpenStudy (big_mensa961):

The slope of the line containing the points (-5, 3) and (-2, 1) is -2/3 2/3 -3/2

Mehek (mehek14):

\(\dfrac{1-3}{-2-(-5)}\)

OpenStudy (big_mensa961):

so i put that in

Mehek (mehek14):

subtract

OpenStudy (skullpatrol):

Simplify it.

OpenStudy (big_mensa961):

1-3=-2/-2-(-5)=3

OpenStudy (big_mensa961):

then wat i Do

Mehek (mehek14):

correct

OpenStudy (skullpatrol):

replace each with their value

Mehek (mehek14):

so what is the numerator and denominator then?

OpenStudy (big_mensa961):

-1=Denominator/ and 3=numerator

Mehek (mehek14):

-1?

OpenStudy (big_mensa961):

-2 i mean

OpenStudy (big_mensa961):

so it -2/3

Mehek (mehek14):

correct

OpenStudy (big_mensa961):

thank you all i have some more taking a test to bring my grade up. from a 70.81/ to a 75